# Learning to Adapt through Bio-Inspired Gait Strategies for Versatile Quadruped Locomotion

### v2

@@ III-C Policy Training
in which $\gamma\in[0,1)$ is the discount factor, $\xi$ is a finite-horizon trajectory dependent on $\pi$ with length $N$, $p(\xi|\pi)$ is the likelihood of $\xi$, and $r$ is the reward function. The proximal policy optimization (PPO) algorithm [51] is used to train all policies and the hyperparameters used are detailed in Supplementary Section 6, which were selecting through the standard method of parameter tuning. As discussed in Section I-A, we estimate the state of the robot during training using an SE. Hence, in terms of applying state feedback noise for domain randomisation to improved sim-to-real transfer we only need to implement this on the input sensor data vector of the SE, $\bm{\sigma}=[\bm{\omega_{B}},\dot{\bm{v}}_{B},\bm{q},\dot{\bm{q}},\bm{\tau},\bm{f}_{\text{grf}}]$. This vector includes base angular velocity, $\bm{\omega_{B}}\in{\mathbb{R}}^{\text{3}}$, base linear acceleration, $\dot{\bm{v}}_{B}\in{\mathbb{R}}^{\text{3}}$, joint positions, $\bm{q}$, joint velocities, $\dot{\bm{q}}$, joint torques, $\bm{\tau}$, and foot ground reaction forces, $\bm{f}_{\text{grf}}\in{\mathbb{R}}^{\text{4}}$. As the initial state of the robot and its performance can never be guaranteed during real-world deployment, we also randomise the initial configuration of the robot, the mass of the robot’s base, $K_{p}$ and $K_{d}$. Additionally, to ensure that a rich variation of $\bm{U}^{\text{cmd}}$ is experienced during training randomly sampled gaits, velocity commands and velocity change durations (to achieve random acceleration) are implemented during training. For all details regarding the noise and sampling used within training please refer to Supplementary Section 7. Although sim-to-real transfer can pose a significant challenge when training DRL policies, we have found that through using domain randomisation, realistic and diverse velocity commands, and generating all robot state observations from the state estimator our framework is able to achieve zero-shot traversal in all experiments and environments shown in Figures 2 and 6, hence demonstrating our methods sufficiently bridge the gap between simulation and the real world. The environment itself is constructed using RaiSim [52], as the vectorized environment setup allows for efficient training of policies. Additionally, the observation normalisation functionality offered by RaiSim is also used for improved training.
@@
During the training of $\pi_{L}$ only flat terrain is present within the environment to isolate and highlight the effect of implementing $\bm{\beta}_{L}$; a core claim of this work is that the implementation of $\bm{\beta}_{L}$ aims to impart gait procedural memory within $\pi_{L}^{\text{bio}}$ hence if rough terrain was observed during training it will become ambiguous if the improved performance is a direct result of implementing $\bm{\beta}_{L}$. However, for training $\pi_{G}$, flat to very rough terrain is implemented using fractal noise, enabling the policy to learn to employ the use of each gait minimising biomechanics metrics on a variety of terrains. We train all variations of $\pi_{L}$ and $\pi_{G}$ for $20$k iterations, taking 6 and 9 hours respectively, on a standard desktop computer with one Nvidia RTX3090 GPU with a training frequency of $100$Hz. It is also important to note that the training of all $\pi_{G}$ policies only utilise our final proposed bio-inspired locomotion framework $\pi_{L}^{\text{bio}}$.
@@ III-E Biomechanics Gait Transition Metrics
As the set of biomechanics metrics applied in this work were originally designed for analysing animal locomotion, several adjustments to how they are calculated needed to be implemented; for example, energy consumption in animals is often measured through the rate of consumption of $\text{O}_{2}$, hence unsuitable for the application of robotics. Additionally, as robots provide a wide array of feedback data some of the metrics have also been augmented to better reflect the characteristic that these biomechanics metrics are attempting to characterise. That being said, for Figure 5(a) only the original biomechanics metrics are applied to allow for direct comparison between robot and animal data.
@@ III-E1 Energy Efficiency
where $P$ is power consumed, $m$ is the system’s mass, and $g$ is gravity. When studying animal locomotion, $P$ is found through measuring how much $\text{CO}_{2}$ is generated and $\text{O}_{2}$ is consumed and $v$ is assumed to be the speed of the treadmill the animal is running on [30, 31, 5]. For the case of the robot, we calculate $P$ from $\bm{\tau}$ and $\dot{\bm{q}}$ with an adjustment term, adopted from [53], and $v$ is assumed to be the magnitude of the robot base velocity command to take a similar approach to animal studies and for consistent metric use between simulation and real-world deplopyment; completely accurate measurement of the robot’s linear base velocity is impossible during real-world deployment due to the accumulation of error within the SE. As such, calculation of the robot’s CoT is formulated as
@@ III-E2 Actuator-structural Forces
As gaining an exact understanding of the actuator-structural forces within animals, researchers have opted instead to measure the peak ground reaction forces of the animal’s stance feet during locomotion using force plates [37]. Other methods include adding strain gauges to the bones of the animals [38]. However, in the case of robots we have the privilege of having access to joint state feedback while also knowing the exact limitations of the hardware. Therefore, when considering the biomechanics hypothesis that animals aim to minimise actuator-structural forces to prevent injury and that torque is proportional to strain and force, in the case of the robot we chose to characterise the actuator-structural forces through joint torque saturation, $\tau_{\%}$, which is calculated by
@@
where $\bm{\tau}_{\text{lim}}\in{\mathbb{R}}^{n}$ is the joint torque limits (assumed based on manufacturers specification), which proves particularly usefully when considering that the hip joints of most quadruped robots, including the A1, are often more sensitive to forces at the foot due to their distance from the point of ground impact and the only motor of the leg in its set alignment; this would not be considered if just ground reaction force was used to characterise actuator-structural forces.
@@ III-E3 Mechanical Work Efficiency
where $t_{\text{gait}}$ is the duration of the current gait cycle, and $\Delta E_{\text{k},i}$ and $\Delta E_{\text{p},i}$ are the changes in kinetic and potential energy over a control time step respectively. The primary difference between the metrics seen in biomechancis to out formulation of $W_{\text{ext}}$ is that $\Delta E_{\text{k},i}$ accounts for not only forward linear velocity but also lateral and angular velocity whereas originally only forward linear velocity is considered.
@@ III-E4 Stability
The best indication of stability in animals is their stride duration CV. This metric characterises periodicity, which is a primary indication of stable locomotion [35]. However, to accurately calculate this, the mean and standard deviation of the stride duration needs to be taken over an extended period of time for appropriate data generation. This is sufficient for undertaking analysis similar to that presented in Figure 5(a), but this presents an issue when it comes to analysing the performance of the proposed control framework as it is common for multiple speed commands being used within the duration of one stride. Hence, to overcome this limitation we instead use $c_{\text{avg}}^{\text{err}}=|\bm{c}^{\text{err}}|/4$, which can be measured every time step rather than just each foot touchdown event; the gait references generated by the BGS have a constant and periodic stride duration, therefore an accurate tracking of this reference would in turn indicate high periodicity, which is further supported by the correlation between the two metrics within Figure 5(b).
@@ III-F Gait Selection Policy
To achieve optimal gait selection for a given state, we leverage the biomechanics metrics within the reward function of $\pi_{G}$, $r_{G}$. For the different variations of $\pi_{G}$ used within Figure 5(a), each policy’s reward function only features the metric that its focusing on within $r_{G}$ but $\pi_{G}^{\text{uni}}$ unifies all metrics hence uses the full form of $r_{G}$ with all metrics. Additionally, as the biomechanics metrics all describe a characteristic that animals try to minimise through changing gaits, they can be directly applied within $r_{G}$ with some normalisation where appropriate. The full form of $r_{G}$ is

### v3

@@ III-C Policy Training
in which $\gamma\in[0,1)$ is the discount factor, $\xi$ is a finite-horizon trajectory dependent on $\pi$ with length $N$, $p(\xi|\pi)$ is the likelihood of $\xi$, and $r$ is the reward function. The proximal policy optimization (PPO) algorithm [52] is used to train all policies and the hyperparameters used are detailed in Supplementary Section 6, which were selecting through the standard method of parameter tuning. As discussed in Section I-A, we estimate the state of the robot during training through using an SE. Hence, in terms of applying state feedback noise for domain randomisation to improved sim-to-real transfer we only need to implement this on the input sensor data vector of the SE, $\bm{\sigma}=[\bm{\omega_{B}},\dot{\bm{v}}_{B},\bm{q},\dot{\bm{q}},\bm{\tau},\bm{f}_{\text{grf}}]$. This vector includes base angular velocity, $\bm{\omega_{B}}\in{\mathbb{R}}^{\text{3}}$, base linear acceleration, $\dot{\bm{v}}_{B}\in{\mathbb{R}}^{\text{3}}$, joint positions, $\bm{q}$, joint velocities, $\dot{\bm{q}}$, joint torques, $\bm{\tau}$, and foot ground reaction forces, $\bm{f}_{\text{grf}}\in{\mathbb{R}}^{\text{4}}$. As the initial state of the robot and its performance can never be guaranteed during real-world deployment, we also randomise the initial configuration of the robot, the mass of the robot’s base, $K_{p}$ and $K_{d}$. Additionally, to ensure that a rich variation of $\bm{U}^{\text{cmd}}$ is experienced during training randomly sampled gaits, velocity commands and velocity change durations (to achieve random acceleration) are implemented. For all details regarding the noise and sampling used within training please refer to Supplementary Section 7. Although sim-to-real transfer can pose a considerable challenge when training DRL policies, we have found that through using domain randomisation, realistic and diverse velocity commands, and generating all robot state observations from the SE, our framework is able to achieve zero-shot traversal in all experiments and environments shown in Figures 2 and 6, hence demonstrating our methods sufficiently bridge the gap between simulation and the real world. The environment itself is constructed using RaiSim [53], as the vectorized environment setup allows for efficient training of policies. Additionally, the observation normalisation functionality offered by RaiSim is also used for improved training.
@@
During the training of $\pi_{L}$ only flat terrain is present within the environment to isolate and highlight the effect of implementing $\bm{\beta}_{L}$. A core claim of this work is that the implementation of $\bm{\beta}_{L}$ aims to impart gait procedural memory within $\pi_{L}^{\text{bio}}$ hence if rough terrain was observed during training it will become ambiguous if the improved performance is a direct result of implementing $\bm{\beta}_{L}$. However, for training $\pi_{G}$, flat to very rough terrain is implemented using fractal noise, enabling the policy to learn to employ the use of each gait minimising biomechanics metrics on a variety of terrains. We train all variations of $\pi_{L}$ and $\pi_{G}$ for $20$k iterations, taking 6 and 9 hours respectively, on a standard desktop computer with one Nvidia RTX3090 GPU with a training frequency of $100$Hz. It is also important to note that the training of all $\pi_{G}$ policies only utilise our final proposed bio-inspired locomotion framework $\pi_{L}^{\text{bio}}$.
@@ III-E Biomechanics Gait Transition Metrics
Although the set of biomechanics metrics applied in this work were designed to accommodate different animals of the same morphology, even when animal body size and weight vary considerably, the fact still remains that they were designed for the analysis of animal locomotion. Hence, several adjustments to how they are calculated needs to be implemented; for example, energy consumption in animals is often measured through the rate of consumption of $\text{O}_{2}$, hence unsuitable for the application of robotics. Additionally, as robots provide a wide array of feedback data some of the metrics have also been augmented to better reflect the characteristic that these biomechanics metrics are attempting to characterise. That being said, for Figure 5a only the original biomechanics metrics are applied to allow for direct comparison between robot and animal data.
@@ III-E1 Energy Efficiency
where $P$ is power consumed and $m$ is the system’s mass. When studying animal locomotion, $P$ is found through measuring how much $\text{CO}_{2}$ is generated and $\text{O}_{2}$ is consumed and $v$ is assumed to be the speed of the treadmill the animal is running on [30, 31, 5]. For the case of the robot, we calculate $P$ from $\bm{\tau}$ and $\dot{\bm{q}}$ with an adjustment term, adopted from [54], and $v$ is assumed to be the magnitude of the robot base velocity command to take a similar approach to animal studies and for consistent metric use between simulation and real-world deplopyment; completely accurate measurement of the robot’s linear base velocity is impossible during real-world deployment due to the accumulation of error within the SE. As such, calculation of the robot’s CoT is formulated as
@@ III-E2 Actuator-structural Forces
As gaining an exact understanding of the actuator-structural forces within animals is infeasible, researchers have opted instead to measure the peak ground reaction forces of the animal’s stance feet during locomotion using force plates [37]. Other methods include adding strain gauges to the bones of the animals [38]. However, in the case of robots we have the privilege of having access to joint state feedback while also knowing the exact limitations of the hardware. Therefore, when considering the biomechanics hypothesis that animals aim to minimise actuator-structural forces to prevent injury and that torque is proportional to strain and force, in the case of the robot we chose to characterise the actuator-structural forces through joint torque saturation, $\tau_{\%}$, which is calculated by
@@
where $\bm{\tau}_{\text{lim}}\in{\mathbb{R}}^{n}$ is the joint torque limits (assumed based on manufacturers specification), which proves particularly usefully when considering that the hip joints of most quadruped robots, including the A1, are often more sensitive to forces at the foot due to their distance from the point of ground impact and the only motor of the leg set in this plane; this would not be considered if just ground reaction force was used to characterise actuator-structural forces.
@@ III-E3 Mechanical Work Efficiency
where $t_{\text{gait}}$ is the duration of the current gait cycle, and $\Delta E_{\text{k},i}$ and $\Delta E_{\text{p},i}$ are the changes in kinetic and potential energy over a control time step respectively. The primary difference between the metrics seen in biomechancis and our formulation of $W_{\text{ext}}$ is that $\Delta E_{\text{k},i}$ accounts for not only forward linear velocity but also lateral and angular velocity whereas originally only forward linear velocity is considered.
@@ III-E4 Stability
The best indication of stability in animals is their stride duration CV. This metric characterises periodicity, which is a primary indication of stable locomotion [35]. However, to accurately calculate this, the mean and standard deviation of the stride duration needs to be taken over an extended period of time for appropriate data generation. This is sufficient for undertaking analysis similar to that presented in Figure 5a, but this presents an issue when it comes to analysing the performance of the proposed control framework as it is common for multiple speed commands being used within the duration of one stride. Hence, to overcome this limitation we instead use $c_{\text{avg}}^{\text{err}}=|\bm{c}^{\text{err}}|/4$, which can be measured every time step rather than just at each foot touchdown event; the gait references generated by the BGS have a constant and periodic stride duration, therefore an accurate tracking of this reference would in turn indicate high periodicity, which is further supported by the correlation between the two metrics within Figure 5b.
@@ III-F Gait Selection Policy
To achieve optimal gait selection for a given state, we leverage the biomechanics metrics within the reward function of $\pi_{G}$, $r_{G}$. For the different variations of $\pi_{G}$ used within Figure 5a, each policy’s reward function only features the metric that its focusing on within $r_{G}$ but $\pi_{G}^{\text{uni}}$ unifies all metrics hence uses the full form of $r_{G}$ with all metrics. In addition, as the biomechanics metrics all describe characteristics that animals try to minimise through changing gaits, they can be directly applied within $r_{G}$ with some normalisation where appropriate. The full form of $r_{G}$ is
@@
Code Availability
Acknowledgments
[…]
Author Contributions Statement
J.H.: conceptualisation, formal analysis, control framework design, software design, simulation design, hardware experiment design, data collection and analysis, wrote the manuscript; C.Z.: conceptulisation, formal analysis, discussion, review and editing, supervision, funding acquisition.